# Three-channel charge Kondo model at high transparency

### v1

@@ Abstract
Quantum impurity models involving a localized charge that is weakly coupled to electronic leads usually map to Kondo-like Hamiltonians that exhibit various quantum critical behaviors. Here, we address the opposite regime of high contact transparency by solving the model of a quantum island coupled to three quantum Hall edge channels. Using a functional renormalization group (FRG) approach, we demonstrate that the low-energy physics is controlled by a nonperturbative fixed point. The universal energy crossover in both the linear conductance and the impurity entropy is obtained. We reproduce the zero-frequency conductance and the leading low-energy exponent of the three-channel Kondo model, confirming their universality across all transparencies. For interacting leads —a model that continuously connects to the pseudo-gap Kondo model at low transparency— we find a line of fixed points as the Luttinger parameter $K$, which encodes the strength of the interactions, changes. Our work demonstrates that FRG methods are an efficient tool for solving quantum impurity problems in regimes where standard approaches fail.
@@ Introduction.
Quantum impurity models, which describe the interaction of a few localized quantum degrees of freedom with a macroscopic bath, provide a framework for exploring various fundamental phenomena, including quantum criticality [1] and exotic states of matter [2], as well as non-Fermi-liquid behavior. Recently, these models have been proposed as potential platforms for hosting localized, fractionalized excitations with anyonic braiding properties [3, 4, 5, 6, 7, 8]. The most prominent model in this class is the Kondo model and its multichannel extensions [9, 10, 11, 12, 13, 14]. This model has been subjected to intense experimental scrutiny [15, 16, 17, 18, 19, 20, 21, 22] in the last three decades, largely driven by the continuous progress in isolating and manipulating single impurities in mesoscopic devices [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38].
Beyond the conventional spin Kondo effect [39], engineering of metallic islands coupled to quantum Hall edge channels enables the implementation of the charge version of the Kondo model [40, 41] with unprecedented control over the experimental parameters. In this setting, the impurity spin is replaced by a pseudospin formed by two competing charge states of the metallic island. Each electron that tunnels between the edge channels and the metallic island flips the pseudospin and a mapping to the Kondo Hamiltonian can be formulated [42] provided that the transparency of the contacts is not too high. In this low-transparency regime, numerical renormalization group (NRG) predictions [43, 44] are very accurate, closely reproducing the experimental data with one [4, 45, 46, 47] or two islands [48, 49].
However, the mapping to the Kondo model breaks down when the contacts are highly transparent and the occupancy of many charge states becomes significant. Nevertheless, the low-energy properties are expected to be the same for all transparencies, even in the absence of an underlying Kondo mapping. The universality of the low-energy behavior has been theoretically demonstrated in the two-channel Kondo (2CK) model [50, 51, 52], and is supported by experimental evidence [46] in the three-channel Kondo (3CK) model as well as by theoretical insights [53, 54, 55]. This is illustrated in Fig. 1 by a sketch of the linear conductance for both low and high transparencies, as observed experimentally in Ref. [46]. In both cases, the same experimental value of the zero-frequency conductance was obtained, in good agreement with the universal value (in units of $e^{2}/h$) $G_{\rm 3CK}^{*}=2\sin^{2}(\pi/5)$ predicted by conformal field theory (CFT) [56, 53, 54, 55].
@@
Figure 1: (Left) Sketch of the experimental device used in [46]. A central metallic island (the Kondo impurity) with a gate-induced offset charge $N_{0}$ is connected to three chiral channels via quantum point contacts (QPC). (Right) Sketch of the conductance $G$ through the island as a function of the energy scale $E$ (temperature or frequency). The red point shows the exact value of the zero-frequency conductance obtained from CFT [56, 53, 54, 55]. The low-transparency (blue) and high-transparency (green) curves schematically represent the measurements performed in Ref. [46] (r=r={}reflection coefficient, t=t={}transmission coefficient). The low-transparency experimental data are reproduced using NRG, while the present work focuses on the high-transparency limit, addressed with the FRG approach.
In this Letter, we explore the high-transparency regime [57] where NRG techniques fail due to the large number of charge states on the island. The three-channel model maps onto a $(0+1)$-dimensional sine-Gordon-like model [53, 54, 56], which we solve using the nonperturbative functional renormalization group (FRG) [58, 59, 60]. We obtain the universal energy crossover in the conductance and the impurity residual entropy as functions of frequency or temperature. This crossover is governed by a single energy scale, $T^{*}$, which depends on microscopic details of the model. The universal value $G^{*}_{\rm 3CK}$ of the zero-frequency conductance is recovered with a 2% accuracy.
The charge 3CK model can be extended to include interactions in the leads. In that case, the one-dimensional edge channels can be viewed as Luttinger liquids characterized by a Luttinger parameter $K\neq 1$. In the low-transparency limit, the model continuously connects to the pseudo-gap Kondo model [61, 62]. FRG calculations reveal a continuous line of nonperturbative low-energy fixed points, characterized by a universal zero-frequency conductance $G^{*}(K)$ and a zero-temperature impurity entropy $\Delta S^{*}(K)$, as $K$ varies. The fully transparent limit, where $G^{*}=e^{2}K/h$, is reached at $K=3/2$, while the zero-transmission limit is reached at $K=2/3$.
@@
We consider the three-channel charge Kondo geometry depicted in Fig. 1 and realized in the experiments of Refs. [45, 46]. When the level spacing in the island is the smallest energy scale in the system, one can neglect the finite volume of the island. Under these conditions, the island and the three channels can be described by a one-dimensional model [40, 41] with the Hamiltonian
@@
where $\hat{\psi}_{j,R/L}$ is the annihilation operator for right/left moving (spinless) electrons in the channel $j$, and $v_{F}$ is the Fermi velocity (we set $\hbar=k_{B}=1$ throughout). The region $x>0$ corresponds to three pairs of counter-propagating spin-polarized quantum Hall edge states, while the region $x<0$ represents the electronic states within the metallic island. Using standard bosonization techniques [63, 64, 65], this Hamiltonian can be written in terms of bosonic operators,
@@
where $N_{0}$ is the gate-induced offset charge and $E_{C}=e^{2}/2C$ with $C$ the effective capacitance of the island. The last contribution to the Hamiltonian accounts for the electron backscattering amplitude $\tilde{r}_{j}$ at each quantum point contact (QPC) between the island and the leads,
@@
Low transparencies are obtained by considering large reflection amplitudes $\tilde{r}_{j}\gg 1$. In this limit, the fields $\hat{\phi}_{j}(0)$ are pinned to zero modulo $\pi$, which corresponds to integer values of the charge $\hat{N}$ on the island. When $N_{0}=1/2$, the charging energy, Eq. (3), is minimized for the two charge states $N=0$ and $N=1$, which form a pseudospin-$1/2$. At temperatures $T\ll E_{C}$, the other charge states can be ignored and the Hamiltonian maps to a Kondo model with three non-interacting conduction channels for $K=1$. As the temperature decreases, the Kondo interaction progressively screens the pseudospin. Initially suppressed due to the low transparency of the contacts, the conductance through the island increases as the temperature is lowered, approaching the zero-temperature 3CK universal value $G^{*}_{3\mathrm{CK}}=2\sin^{2}(\pi/5)$ [53], which reflects partial screening of the impurity, see Fig. 1. This result is consistent with experimental observations [46].
Large transparencies, on the other hand, correspond to small reflection amplitudes $\tilde{r}_{j}\ll 1$. In this regime, it is convenient to integrate out the bosonic fields away from the contact ($x\neq 0$) along with the gapped charge mode [66], resulting in the Euclidean action (for $N_{0}=1/2$)
@@
where ${\bm{\varphi}}(\tau)$ is a two-component vector that depends linearly on the original three fields $\phi_{1,2,3}$. The imaginary time $\tau\in[0,\beta]$, with $\beta=1/T$, and $\omega_{n}=2n\pi/\beta$ ($n\in\mathbb{Z}$) is a Matsubara frequency. The periodic potential in (5) exhibits the symmetry of a triangular lattice defined by the reciprocal vectors ${\bm{G}}_{1}=(1,-1/\sqrt{3})$ and ${\bm{G}}_{2}=(0,2/\sqrt{3})$, with ${\bm{G}}_{3}=-{\bm{G}}_{1}-{\bm{G}}_{2}$. The effective reflection coefficients are defined by
@@
where $\gamma$ is the Euler constant [66]. In the following, we focus on the symmetric case where all reflection coefficients are equal, $r_{j}\equiv r$. The action (5), supplemented by the ultraviolet (UV) energy cutoff $D$, also describes the quantum Brownian motion of a particle with coordinate $\bm{\varphi}$} in a two-dimensional periodic potential [53, 54, 56].
@@
The reflection coefficient $r$ has a scaling dimension 12K/31-2K/3 and is a relevant perturbation for $K<3/2$. A perturbative treatment is justified near $K=3/2$, but a nonperturbative approach is required elsewhere. Here, we use the nonperturbative FRG approach [67], which is a modern implementation of Wilson’s RG. This approach enables us to calculate the effective action (or Gibbs free energy in statistical physics terminology) from an exact functional flow equation obtained by progressively integrating out high-energy degrees of freedom [66]. The effective action
@@
Figure 2 shows the zero-temperature conductance for non-interacting leads ($K=1$). As expected for an island in the high-transparency regime, the conductance approaches unity at high frequencies. At low frequencies, it decreases significantly, indicating a reduction in the effective transparency of the QPCs. While the conductance typically depends on the UV energy cutoff $D$ and the reflection coefficient $r$, in the wide-band limit it exhibits universal scaling as a function of the rescaled frequency $\omega/T^{*}$,
@@
$\frac{T^{*}}{D}\propto\left(\frac{r}{D}\right)^{3}.$ (11)
This is in perfect agreement with the FRG results [66]. In practice, we define $T^{*}$ such that $G(\omega=T^{*},T=0)$ lies halfway between $1$ and the zero-energy conductance $G^{*}_{\rm 3CK}=G(\omega=0,T=0)=\mathcal{G}(0)$.
Figure 2: (Left) Conductance (in units of $e^{2}/h$) as a function of rescaled frequency $\omega/T^{*}$ for various values of $r$ with $K=1$. The asymptotic behavior at low-frequency (13) and high-frequency (12) is shown by (red) dashed lines, while the horizontal dash-dotted line marks the extrapolated zero-frequency limit. The red point indicates the exact result at zero frequency obtained from CFT. The curves collapse onto a single line (the black solid line is a guide to the eye) in the wide-band limit $T^{*}/D\ll 1$. (Right) Universal entropy crossover for $K=1$, compared with the exact thermodynamic Bethe ansatz (TBA) solution.
@@
where $G^{*}_{\rm 3CK}=2\sin^{2}(\pi/5)$ and $\nu=2/5$ is the exponent of the leading irrelevant operator. The FRG yields very accurate estimates of $G^{*}_{\rm 3CK}$ and $\nu$, as summarized in Table 1. In addition to providing a benchmark for assessing the accuracy of the FRG, the case $K=1$ also illustrates the ability of the method to access real-frequency dynamical quantities such as $G(\omega)$, which are difficult to compute with other techniques.
The residual impurity entropy exhibits a similar scaling form in the large bandwidth limit $T^{*}\ll D$,
@@
with $\mathcal{S}$ a universal function. At high temperatures, transport through the QPCs is nearly ballistic, and the residual entropy remains close to zero. It decreases with temperature and takes negative values, larger than $\ln 2$, indicating that the impurity is partially screened (complete screening would give $\Delta S(0)=-\ln 2\simeq-0.693$). Whereas the conductance is known exactly only at zero temperature and frequency, the full temperature dependence of the entropy can be obtained from the thermodynamic Bethe ansatz (TBA) equations [56, 66]. The universal scaling function obtained from FRG is nearly identical to the exact result over a broad temperature range $T/T^{*}\gtrsim 10^{-1}$. Eventually, it deviates from the exact TBA solution in the low-temperature limit, although $\mathcal{S}(0)$ remains within 9% of the exact value. The exact result, $\mathcal{S}(0)=\ln[(1+\sqrt{5})/2\sqrt{3}]$, has been interpreted as a signature of emerging Fibonacci anyons [4, 5]. Whether the FRG captures these (putative) low-energy excitations remains an open question.
@@
Figure 3: Zero-frequency conductance $G^{*}(K)$ as a function of $K$, along with the corresponding exponent $\nu(K)$ (right inset). The dashed lines indicate perturbative RG results near $K=3/2$, and near $K=2/3$ obtained from the dual model [53, 54]. The exponent $\nu$ reaches its maximum around $K=1$ (right inset). The left inset shows the residual impurity entropy $\Delta S^{*}(K)$. Solid (black) lines are guides to the eye.
Electron interactions in the leads can be taken into account by introducing a Luttinger parameter $K$ different from 1. In this case, no exact results are available from CFT or TBA. However, a phase diagram was conjectured based on a perturbative RG analysis [53, 54]. Depending on the value of the Luttinger parameter, three distinct regions are expected. For $K>3/2$, the weak-coupling fixed point $r=0$ is stable, and the conductance approaches 1 at zero temperature and frequency. For $K<2/3$, the strong-coupling fixed point, which can be studied perturbatively in the dual model [53, 54], is stable. This causes the conductance to vanish. In the intermediate range $2/3<K<3/2$, neither the weak-coupling fixed point nor the strong-coupling fixed points are stable. Instead, we expect a line of intermediate stable fixed points that exhibit Kondo-type physics, as in the case $K=1$.
Our analysis fully endorses this scenario. The FRG reproduces the ballistic conductance $G(\omega=0,T=0)=1$ for $K>3/2$ and reveals the line of intermediate fixed points that emerges as $K$ decreases. The perturbative RG results near $K=3/2$ from Ref. [54] are recovered. The behavior for $2/3<K<3/2$ parallels that at $K=1$: The conductance approaches $1$ at high frequencies and a nonzero value $G^{*}(K)<1$ in the zero-frequency limit (Fig. 3). The residual entropy remains close to zero at high temperatures and reaches a finite value $\Delta S^{*}(K)$ at $T=0$ (Fig. 2). The universal scaling forms  (10) and (14) of conductance and residual entropy extend beyond the case $K=1$ and remain valid throughout the intermediate region (see Fig. 4). Thus, the FRG provides quantitative predictions for the conductance and residual entropy over a wide range of $K$ values, which could be tested in future experiments. Note that the range of $K$ we can study is limited, as integrating the FRG equations becomes increasingly computationally expensive for $K\lesssim 0.9$.
Figure 4: Universal conductance scaling function $\mathcal{G}$ for various values of $K$. The markers correspond to the ratios $r/D\in\{0.005,0.01,0.05\}$. Lines are guides to the eye.

### v2

# Universal crossover in the three-channel charge Kondo model at high transparency

@@ Abstract
Quantum impurity models provide a central framework for correlated electron physics, with quantum dots enabling controlled experimental realizations. While their weak-coupling behavior is well understood through mappings to Kondo Hamiltonians, the opposite regime of highly transparent contacts has lacked a controlled theoretical description. Using the functional renormalization group (FRG), we resolve this regime for the three-channel charge Kondo device of Ref. [1], benchmarking against conformal field theory by reproducing the universal zero-frequency conductance and, crucially, going beyond it to obtain the full frequency crossover of the conductance and the full temperature crossover of the impurity entropy, together with a continuous line of fixed points for interacting leads. These results establish FRG as a powerful nonperturbative tool for quantum impurity problems in regimes inaccessible to conventional approaches, with direct implications for mesoscopic experiments.
@@ Introduction.
Quantum impurity physics provides a window into how a small quantum system can dramatically influence a much larger environment. The classic example is the Kondo effect, where a single magnetic impurity in a metal produces a collective many-body state that reshapes electronic transport [2, 3]. Multichannel extensions reveal exotic non-Fermi-liquid behavior and quantum criticality that have become cornerstones of modern condensed matter theory [4, 5, 6, 7]. Beyond their conceptual importance, impurity models have become central to nanoscale experiments across a variety of platforms, ranging from scanning tunneling spectroscopy on magnetic adatoms [8, 9, 10, 11, 12, 13, 14] to semiconductor quantum dots [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], which can be tuned with unprecedented control. These developments have opened the door to realizing exotic correlated states with potential connections to fractionalized excitations and anyonic physics [32, 33, 34, 35, 36, 37]. A central challenge is to understand how universality survives when the standard mapping to Kondo Hamiltonians breaks down.
Metallic islands coupled to quantum Hall edge channels provide an ideal platform for exploring this problem. In the low-transparency regime, tunneling involves only two charge states, and the system maps onto charge versions of multichannel Kondo models [38, 39, 40]. In this regime, numerical renormalization group (NRG) calculations are quantitatively accurate [41, 42] and have been confirmed experimentally in one- and two-island geometries [33, 43, 1, 44, 45, 46]. A hallmark of this regime is the universality of the low-energy behavior, established in the two-channel Kondo model through theory [47, 48, 49] and experiment [1]. However, at high contact transparency the mapping fails, as many charge states contribute simultaneously. Despite this breakdown, experiments on the three-channel case suggest that universality persists, with the same zero-frequency conductance observed in both limits [1], as illustrated in Fig. 1.
@@
Figure 1: (Left) Experimental device from Ref. [1]: a metallic island (the Kondo impurity) with gate-induced offset charge $N_{0}$ is coupled to three quantum Hall edge channels via quantum point contacts. (Right) Schematic conductance $G$ as a function of energy scale $E$ (temperature or frequency) for low (blue) and high (green) transparencies. Both regimes yield the same universal zero-frequency conductance $G_{\rm 3CK}^{*}=2\sin^{2}(\pi/5)$ (red point), as predicted by conformal field theory [50, 51, 52, 53].
In this Letter, we provide a theoretical framework for this strongly coupled regime. Focusing on the charge three-channel Kondo model (3CK), we show that its effective description maps onto a $(0+1)$-dimensional sine-Gordon-like model [51, 52, 50], which we solve using a nonperturbative functional renormalization group (FRG) approach [54, 55, 56]. We obtain the full universal crossover of the conductance and impurity entropy, controlled by a single energy scale $T^{*}$. The corresponding universal scaling functions are precisely the quantities probed experimentally, where data are collapsed as functions of frequency or temperature scaled by $T^{*}$ and compared to theoretical curves. Our results reproduce the exact zero-frequency conductance of conformal field theory (CFT) with 2% accuracy, establishing the universality of the low-energy physics across all transparencies [51, 52, 53]. Extending the analysis to interacting leads, we reveal a continuous line of nonperturbative fixed points, parameterized by the Luttinger parameter $K$ and connected to pseudo-gap Kondo physics [57, 58]. These findings demonstrate that FRG can access impurity regimes beyond the reach of standard methods, providing quantitative predictions for ongoing experiments.
@@
We consider the three-channel charge Kondo geometry depicted in Fig. 1 and realized experimentally in Refs. [43, 1]. When the single-particle level spacing in the island is the smallest energy scale in the system, the island can be treated as a zero-dimensional object. In this limit, the island and the three leads are described by a one-dimensional model [39, 40] with Hamiltonian
@@
where $\hat{\psi}_{j,R/L}$ annihilates a right/left-moving spinless electron in channel $j$, and $v_{F}$ is the Fermi velocity (we set $\hbar=k_{B}=1$). The region $x>0$ corresponds to the three pairs of counter-propagating spin-polarized quantum Hall edge states, while the region $x<0$ represents electronic states inside the metallic island. Using standard bosonization techniques [59, 60, 61], Eq. (1) becomes
@@
where $N_{0}$ is the gate-induced offset charge and $E_{C}=e^{2}/2C$ with $C$ the island effective capacitance. The last contribution to the Hamiltonian accounts for the electron backscattering amplitude $\tilde{r}_{j}$ at each quantum point contact (QPC) between the island and the leads,
$\hat{H}_{\rm QPC}=v_{F}\sum_{j=1}^{3}\tilde{r}_{j}(\hat{\psi}^{\dagger}_{j,R}(0)\hat{\psi}_{j,L}(0)+{\rm H.c.}).$ (4)
In bosonized form, this term reads [62]
@@
For low transparencies ($\tilde{r}_{j}\gg 1$), the fields $\hat{\phi}_{j}(0)$ are pinned to zero modulo $\pi$, corresponding to integer island charge. At $N_{0}=1/2$, the charging energy [Eq. (3)] selects the two charge states $N=0,1$, which form a pseudospin-$1/2$. At temperatures $T\ll E_{C}$, the other charge states can be ignored and the Hamiltonian maps to a Kondo model with three non-interacting conduction channels for $K=1$. As the temperature decreases, the Kondo interaction progressively screens the pseudospin. Initially suppressed due to the low transparency of the contacts, the conductance through the island increases as the temperature is lowered, approaching the zero-temperature 3CK universal value $G^{*}_{3\mathrm{CK}}=2\sin^{2}(\pi/5)$ [51], which reflects partial screening of the impurity, see Fig. 1. This result is consistent with experimental observations [1].
In the opposite limit of high transparency ($\tilde{r}_{j}\ll 1$), many charge states contribute simultaneously. Integrating out the bosonic fields away from the contacts ($x\neq 0$) and the gapped charge mode [62] yields the Euclidean action (for $N_{0}=1/2$),
@@
where ${\bm{\varphi}}(\tau)$ is a two-component field linearly related to $\phi_{1,2,3}$. The imaginary time $\tau\in[0,\beta]$, with $\beta=1/T$, and $\omega_{n}=2n\pi/\beta$ ($n\in\mathbb{Z}$) is a Matsubara frequency. The reciprocal lattice vectors ${\bm{G}}_{1}=(1,-1/\sqrt{3})$, ${\bm{G}}_{2}=(0,2/\sqrt{3})$, and ${\bm{G}}_{3}=-{\bm{G}}_{1}-{\bm{G}}_{2}$ define a triangular potential landscape. The effective reflection coefficients are
@@
with $\gamma$ the Euler constant [62]. In the following, we focus on the symmetric case where all reflection coefficients are equal, $r_{j}\equiv r$. The action (S15), supplemented by the ultraviolet (UV) energy cutoff $D$, also describes the quantum Brownian motion of a particle with coordinate $\bm{\varphi}$} in a two-dimensional periodic potential [51, 52, 50]. Although Eq. (S15) is not, strictly speaking, a Kondo model, its low-energy fixed point, which is discussed below, coincides with that of the three-channel charge Kondo model  [^1] In the conventional charge Kondo formulation (low-transparency regime), the renormalization-group flow from the zero-transparency regime toward the 3CK fixed point takes place in the $SU(2)_{3}$ sector. In contrast, the flow from the fully-transparent regime to the 3CK fixed point occurs in the $SU(2)_{3}/U(1)$ sector [50, 53]..
@@
The reflection coefficient $r$ has a scaling dimension 12K/31-2K/3, hence it is relevant for $K<3/2$. While perturbation theory applies near $K=3/2$, a nonperturbative treatment is required otherwise. We employ the nonperturbative FRG [64], a modern formulation of Wilson’s RG, which provides the effective action $\Gamma[\bm{\phi}]$ (or Gibbs free energy in statistical physics terminology) from an exact functional flow equation obtained by progressively integrating out high-energy degrees of freedom [62]. The effective action
@@
For noninteracting leads, the zero-temperature conductance approaches unity at high frequency and decreases strongly at low frequency, reflecting the renormalization of QPC transparency (Fig. 2). While the conductance typically depends on the UV energy cutoff $D$ and the reflection coefficient $r$, in the wide-band limit it exhibits universal scaling as a function of the rescaled frequency $\omega/T^{*}$,
@@
$\frac{T^{*}}{D}\propto\left(\frac{r}{D}\right)^{3},$ (12)
in agreement with FRG [62]. In practice, we define $T^{*}$ such that $G(\omega=T^{*},T=0)$ lies halfway between $1$ and the zero-energy conductance $G^{*}_{\rm 3CK}=G(\omega=0,T=0)=\mathcal{G}(0)$.
Figure 2: (Left) Universal conductance (in units of $e^{2}/h$) as a function of rescaled frequency $\omega/T^{*}$ for several $r$ values at $K=1$. The asymptotic behavior at low-frequency (14) and high-frequency (13) is shown by (red) dashed lines, while the horizontal dash-dotted line marks the extrapolated zero-frequency limit. The red point marks the exact CFT result. Curves collapse onto a single line (the black solid line is a guide to the eye) in the wide-band limit $T^{*}/D\ll 1$. (Right) Universal entropy crossover at $K=1$, compared with exact thermodynamic Bethe ansatz (TBA) results.
@@
where $G^{*}_{\rm 3CK}=2\sin^{2}(\pi/5)$ and $\nu=2/5$ is the exponent of the leading irrelevant operator. The FRG reproduces both values with remarkable precision (Table 1). In addition to providing a benchmark for assessing the accuracy of our method, the case $K=1$ also illustrates its ability to access real-frequency dynamical quantities such as $G(\omega)$, which are difficult to compute with other techniques.
The residual impurity entropy exhibits similar scaling in the large bandwidth limit $T^{*}\ll D$,
@@
with $\mathcal{S}$ a universal function. At high temperatures, transport through the QPCs is nearly ballistic, and the residual entropy remains close to zero. It decreases with temperature and takes negative values, larger than $\ln 2$, indicating that the impurity is partially screened (complete screening would give $\Delta S(0)=-\ln 2\simeq-0.693$). Whereas the conductance is known exactly only at zero temperature and frequency, the full temperature dependence of the entropy can be obtained from the thermodynamic Bethe ansatz (TBA) equations [50, 62]. The universal scaling function obtained from FRG is nearly identical to the exact result over a broad temperature range $T/T^{*}\gtrsim 10^{-1}$. Eventually, it deviates from the exact TBA solution in the low-temperature limit, although $\mathcal{S}(0)$ remains within 9% of the exact value. The exact result, $\mathcal{S}(0)=\ln[(1+\sqrt{5})/2\sqrt{3}]$, has been interpreted as a signature of emerging Fibonacci anyons [65, 66, 33, 34]. Whether the FRG captures these (putative) low-energy excitations remains an open question.
@@
Figure 3: Zero-frequency conductance $G^{*}(K)$ as a function of $K$, along with the corresponding exponent $\nu(K)$ (right inset). Dashed lines indicate perturbative RG results near $K=3/2$, and near $K=2/3$ obtained from the dual model [51, 52]. The exponent $\nu$ peaks around $K=1$ (right inset). The left inset shows the residual impurity entropy $\Delta S^{*}(K)$. Solid (black) lines are guides to the eye.
Electron interactions in the leads are accounted for by a Luttinger parameter $K\neq 1$. In this regime, exact CFT or TBA solutions are unavailable, but perturbative RG analyses [51, 52] predict a rich phase diagram with three distinct regions. For $K>3/2$, the weak-coupling fixed point ($r=0$) is stable, yielding unit conductance at zero temperature and frequency. For $K<2/3$, the strong-coupling fixed point, which can be studied perturbatively in the dual model [51, 52], is stable and the conductance vanishes. Between these limits, $2/3<K<3/2$, both extremes are unstable, giving rise to a continuous line of intermediate fixed points displaying Kondo-type physics, as in the case $K=1$.
Our FRG analysis fully endorses this scenario. The FRG reproduces the ballistic conductance $G(\omega=0,T=0)=1$ for $K>3/2$ and reveals the line of intermediate fixed points that emerges as $K$ decreases. The perturbative RG results near $K=3/2$ from Ref. [52] are recovered. The behavior for $2/3<K<3/2$ parallels that at $K=1$: The conductance approaches $1$ at high frequencies and a nonzero value $G^{*}(K)<1$ in the zero-frequency limit (Fig. 3). The residual entropy remains close to zero at high temperatures and reaches a finite value $\Delta S^{*}(K)$ at $T=0$ (Fig. 3). The universal scaling forms  (11) and (15) of conductance and residual entropy extend beyond the case $K=1$ and remain valid throughout the intermediate region (see Fig. 4). Thus, the FRG provides quantitative predictions for the conductance and residual entropy over a wide range of $K$ values, which could be tested in future experiments. Note that the range of $K$ we can study is limited, as integrating the FRG equations becomes increasingly computationally expensive for $K\lesssim 0.9$.
Figure 4: Universal conductance scaling function $\mathcal{G}$ for several $K$ values. Markers correspond to $r/D\in\{0.005,0.01,0.05\}$. Lines are guides to the eye.